6. Find the projection of u onto vifu= (-1, - 4), v=(-3,5). ( )
Answer #1

Given U= (-1,-43 H = 11ull = litle = 1 3,57 llull= √9+25= . The projection of uon IV 2 wor= [)(-3) + 1-4165) = 3 – 20 = -17

Given Options are wrong